# Escaping the Verifier: Learning to Reason via Demonstrations

### v1

@@ 3.4 RARO: Relativistic Adversarial Reasoning Optimization
To ensure stable and efficient learning, we implement several optimizations. First, we use a shared LLM for both the critic and the policy, which reduces memory usage and promotes generalization. This allows us to employ data mixing, where policy and critic rollouts are combined in a single batch, simplifying the training loop. To prevent the critic from suffering from catastrophic forgetting, we utilize a replay buffer that mixes past policy rollouts with current ones. Finally, we incorporate several practical improvements to the GRPO algorithm, such as over-length filtering and removing advantage/length normalization. For full implementation details, please refer to Appendix C.1.
@@
Inputs: Dataset $D=\{(q_{i},a_{i})\}$; Tie reward $\tau_{\mathrm{pol}},\tau_{\mathrm{crit}}$; Loss weight $\lambda_{\mathrm{pol}},\lambda_{\mathrm{crit}}$ Batch $B$; Rollout $K$.
@@ 4.1 Tasks & Datasets
Then, we evaluate our method on the domain of general math reasoning problems using the DeepMath dataset (he2025deepmath103klargescalechallengingdecontaminated). Compared to Countdown, answer verification in the general math domain is significantly more challenging, often requiring reproduction of the derivation. Through this task, we aim to stress test our method on difficult general reasoning environments where verification is as difficult as generation.
@@
Finally, we extend our method to its intended setting of non-verifiable, open-ended reasoning tasks using a custom Poetry Writing dataset. Unlike the math tasks, poetry writing does not admit an objective verifier. Thus, for evaluation, we use GPT-5  (openai2025gpt5) as a judge to evaluate poems in both isolation and in comparison to the expert poem (see Appendix C.2 for details). This task represents the non-verifiable regime that our method aims to handle, where explicit reasoning could significantly improve quality.
@@ 4.2 Baselines
The SFT baseline trains the base models to directly maximizes the conditional log-likelihood of the expert answer given the question, representing the standard use of demonstration data.
@@ 4.3 Training & Evaluation Setup
We evaluate our method and baselines on the Qwen2.5 (qwen2025qwen25technicalreport) family of models, and to focus on improving reasoning performance rather than language understanding, we initialize from the instruction-tuned checkpoints instead of the pretrained model checkpoints. We select the Qwen2.5 family they are popular non-reasoning LLMs, allowing us to study the effectiveness of our method on eliciting reasoning behaviors in a controlled manner.
Countdown and DeepMath are evaluated with a ground-truth verifier, while Poetry Writing is evaluated with GPT-5 as a judge in two fashions: a scalar score normalized to 0-100 and a win-rate against the expert poem. See Appendix C.2 for further details.
Each dataset is split into train, validation, and test sets, and we select our checkpoints based on the highest validation performance. For each dataset and model size, we match dataset splits, rollout budgets, hyperparameters, and sampling configurations when possible to ensure a fair comparison. Unless otherwise specified, all methods are trained and evaluated with a reasoning budget of 2048 tokens. Full implementation details and hyperparameters are provided in Appendix C.
4.4 Main Results
@@ 4.4.1 Countdown
accuracy ($\%$) $\uparrow$ score (0-100) $\uparrow$ win-rate ($\%$) $\uparrow$
@@
4.4.2 DeepMath
Next, we evaluate RARO on the DeepMath dataset, a collection of general math problems. For the DeepMath task, we focus on scaling our method and baselines with respect to model size instead of reasoning budget, as it is a much more general setting where model capacity is a real bottleneck in performance.
@@ 4.4.2 DeepMath
As reported in Table 2, RARO consistently outperforms all verifier-free baselines across model scales. With the 1.5B model, we achieve $41.3\%$ accuracy compared to $37.7\%$ for the best baseline (RL-Logit), an improvement of $3.6\%$. This advantage grows with model size: at 3B, our method ($49.1\%$) surpasses the best baseline (RL-Logit, $43.1\%$) by $6.0\%$, and at 7B, it reaches $57.5\%$, beating the best baseline (RL-Logit, $49.3\%$) by $8.2\%$. These results demonstrate that our adversarial learning framework provides a strong signal for reasoning that outperforms not only purely supervised approaches like SFT or Rationalization but also RL-based approaches like RL-Logit.
@@
We further analyze the training dynamics of RARO. As shown in Figure 4 and Figure 7, our coupled training objective maintains a robust equilibrium, allowing the policy to steadily improve its reasoning capabilities and response length without collapsing. This stability confirms the robustness of our optimization procedure.
@@
Another key advantage of RARO is that our learned critic enables natural Test-Time Scaling (TTS) to further improve the policy’s performance. Specifically, our critic’s pariwise comparison setup allows for a single-elimination tournament with the critic as the judge (see Algorithm 3), enabling further policy improvements with additional rollouts. As shown in Figure 2 (and detailed in Table 9), increasing the number of rollouts from 1 to 16 consistently improves performance. Notably, with 16 rollouts, RARO achieves $53.6\%$ on the 1.5B model and $57.9\%$ on the 3B model. When comparing against the RLVR baseline with the same TTS strategy (Table 9), we observe that RARO achieves a similar rate of improvement. This result highlights that RARO, when combined with test-time search, can scale effectively, matching the scaling trends of models trained with oracle verifiers.
4.4.3 Poetry Writing
@@
Table 12 reveals a striking performance gap between RARO and baselines. While SFT and Rationalization achieve modest win-rates against expert poetry (peaking at $5.9\%$ with the 7B model), RARO reaches $25.0\%$, a four-fold improvement. This advantage is also reflected in the scoring evaluation, where RARO consistently surpasses baselines (e.g., $67.8$ vs. $53.7$ for SFT at 1.5B). Notably, RL-Logit, leading baseline for DeepMath, fails to produce competitive results, yielding near-zero improvement over the base model ($0.364$ vs. $0.350$ at 1.5B) for both the win-rate and scoring evaluation. This suggests that standard preference optimization methods and RL-based approaches struggle to generalize in open-ended domains.
@@
A key result is the scalability of RARO with model size in the creative domain. As we increase model capacity from 1.5B to 7B, the win-rate against expert human poems grows substantially, from $7.8\%$ to $25.0\%$. The scoring evaluation similarly improves from $67.8$ to $77.3$. This trend shows that just like verifiable domains, RARO continues to effective scale with model size in open-ended domains.
@@
Qualitatively, RARO induces explicit planning and reasoning behaviors even in open-ended domains. As shown in Figure 5 and fully detailed in Figure 17, the model learns to decompose the prompt into key themes (e.g., “disillusionment”, “transience of power”) and stylistic constraints (e.g., “flowing, rhythmic yet contemplative style”) before generating the poem. This demonstrate that RARO effectively elicits reasonings that align the model’s output to creative poems while adhering to the prompt’s nuanced requirements.
Countdown Input Use the four numbers 5, 33, 40, and 26 to create an expression that equals 24.
Countdown Reasoning & Answer To solve this problem, we need to use the numbers 5, 33, 40, and 26 to create an expression that equals 24. Let’s explore different combinations and find one that meets the requirements. 1. (40 - 26) * (33 - 5) = 14 * 28 = 392 (too high) 2. (40 - 5) * (33 - 26) = 35 * 7 = 245 (too high) 16. (33 + 5) - (40 - 26) = 38 - 14 = 24 (just right! This is the correct solution.) Answer: (33+5)(4026)
DeepMath Input Find a positive integer $n>1$ that is both a sum of two squares ($n=a^{2}+b^{2}$) and a perfect cube ($n=z^{3}$), where $a\neq b$. Determine the non-trivial solutions to the equation $a^{2}+b^{2}=z^{3}$.
DeepMath Reasoning & Answer To find a positive integer $n>1$ that is both a sum of two squares and a perfect cube, where $a\neq b$, we start by considering the general form of such numbers. For $z=2$, we need $n=2^{3}=8$. We check if 8 can be expressed as $a^{2}+b^{2}$ with $a\neq b$ For $z=5$, we need $n=5^{3}=125$. The possible pairs $(a,b)$ are $(3,4)$ and $(4,3)$, and since $a\neq b$, this pair works. Therefore, $125=3^{2}+4^{2}$. Answer: 125
Poetry Input A short free-verse poem about the ironic pageantry of war glory turning to dust at sunset
Poetry Reasoning & Answer To write a free-verse poem about war glory, I should first consider the themes of disillusionment, the transience of power, and what aspects of war can be attributed to pageantry. These should be woven into a flowing, rhythmic yet contemplative style, reflective of an artist’s perspective on war. Answer:
The Decoration (1942)
Nestled, red-glazed in a rail, Is the body of Brigadier Sir Thomas Hassall. His very gold is dust.
Figure 5: Qualitative examples for Countdown, DeepMath, and Poetry Writing at 1.5B. For each task, we show the input (top) and the the truncated reasoning and answer sampled from 1.5B models trained with RARO (bottom). See Appendix E for full examples.
@@ 5 Conclusion & Future Work
We intoduced RARO (Relativistic Adversarial Reasoning Optimization), a novel approach to training reasoning LLMs using only expert demonstrations, thereby bypassing the need for task-specific verifiers or expensive preference annotations. By formulating the problem as Inverse Reinforcement Learning and incorporating a relativistic critic setup, we obtain a principled and stable adversarial training algorithm that yields strong reasoning capabilities.
Our experiments demonstrate the effectiveness of RARO: (i) on the controlled Countdown task, it not only outperforms verifier-free baselines but also nearly matches the performance of RLVR; (ii) on the general math domain, it exhibits similar scalability trends to RLVR while outperforming baselines without verification; and (iii) on the open-ended Poetry Writing task, it successfully elicits emerging specialized reasoning capabilities and significantly surpasses all baselines. Together, these findings suggest that RARO is a promising and practical approach for training reasoning models without reliance on explicit verifiers.
@@ A.2 Proof of Reward Gradient
Using the closed-form policy, the gradient of the data log-likelihood $\mathcal{L}(\phi)=\mathbb{E}_{(q,a)\sim\hat{p}_{D}}\!\left[\log\pi_{\theta^{\star}(\phi)}(a\mid q)\right]$ with respect to $\phi$ is:
@@ A.3 Derivation of Reasoning Reward Gradient
Under the binary classification parameterization $r_{\phi}(a,q)=c_{\phi}(\ell=\mathrm{expert}\mid a,q)-c_{\phi}(\ell=\mathrm{policy}\mid a,q)$, the gradient of the loss with respect to critic parameters $\phi$ is:
@@
We parameterize the reward using a binary classifier (critic) $c_{\phi}(\ell\mid a,q)$ where $\ell\in\{\mathrm{expert},\mathrm{policy}\}$:
$r_{\phi}(a,q)=c_{\phi}(\ell=\mathrm{expert}\mid a,q)-c_{\phi}(\ell=\mathrm{policy}\mid a,q).$
@@
$\mathbb{E}_{\ell\sim c_{\phi}(\cdot\mid a,q)}[\tilde{R}(\ell)\nabla_{\phi}\log c_{\phi}(\ell\mid a,q)]$ $=\tilde{R}(\mathrm{expert})p_{E}\nabla_{\phi}\log p_{E}+\tilde{R}(\mathrm{policy})p_{P}\nabla_{\phi}\log p_{P}$
$=\tilde{R}(\mathrm{expert})\nabla_{\phi}p_{E}+\tilde{R}(\mathrm{policy})\nabla_{\phi}p_{P}$
$=\tilde{R}(\mathrm{expert})\nabla_{\phi}p_{E}+\tilde{R}(\mathrm{policy})(-\nabla_{\phi}p_{E})$
$=(\tilde{R}(\mathrm{expert})-\tilde{R}(\mathrm{policy}))\nabla_{\phi}p_{E}.$
By setting $\tilde{R}(\mathrm{expert})=1$ and $\tilde{R}(\mathrm{policy})=-1$, we obtain:
$(1-(-1))\nabla_{\phi}p_{E}=2\nabla_{\phi}p_{E}=\nabla_{\phi}r_{\phi}(a,q).$
@@
$\nabla_{\phi}r_{\phi}(a,q)=\mathbb{E}_{\ell\sim c_{\phi}(\cdot\mid a,q)}\Big[(\mathbbm{1}_{\ell=\mathrm{expert}}-\mathbbm{1}_{\ell=\mathrm{policy}})\nabla_{\phi}\log c_{\phi}(\ell\mid a,q)\Big].$
@@
1. Expert Term ($(q,a)\sim\hat{p}_{D}$):
$\mathbb{E}_{a\sim\hat{p}_{a\mid q}}[\nabla_{\phi}r_{\phi}(a,q)]$ $=\mathbb{E}_{a\sim\hat{p}_{a\mid q}}\Big[\mathbb{E}_{\ell\sim c_{\phi}}\Big[(\mathbbm{1}_{\ell=\mathrm{expert}}-\mathbbm{1}_{\ell=\mathrm{policy}})\nabla_{\phi}\log c_{\phi}(\ell\mid a,q)\Big]\Big].$
This corresponds to a reward signal of +1+1 when $\ell=\mathrm{expert}$ and 1-1 when $\ell=\mathrm{policy}$.
@@ 1. Expert Term ($(q,a)\sim\hat{p}_{D}$):
$-\mathbb{E}_{a\sim\pi_{\theta}}[\nabla_{\phi}r_{\phi}(a,q)]$ $=\mathbb{E}_{a\sim\pi_{\theta}}\Big[-\mathbb{E}_{\ell\sim c_{\phi}}\Big[(\mathbbm{1}_{\ell=\mathrm{expert}}-\mathbbm{1}_{\ell=\mathrm{policy}})\nabla_{\phi}\log c_{\phi}(\ell\mid a,q)\Big]\Big]$
$=\mathbb{E}_{a\sim\pi_{\theta}}\Big[\mathbb{E}_{\ell\sim c_{\phi}}\Big[(\mathbbm{1}_{\ell=\mathrm{policy}}-\mathbbm{1}_{\ell=\mathrm{expert}})\nabla_{\phi}\log c_{\phi}(\ell\mid a,q)\Big]\Big].$
This corresponds to a reward signal of 1-1 when $\ell=\mathrm{expert}$ and +1+1 when $\ell=\mathrm{policy}$.
@@
This matches the formulation in the method section. ∎
Proposition A.4.
The estimator $\hat{r}_{\phi}(a,q)=\mathbbm{1}_{\ell=\mathrm{expert}}-\mathbbm{1}_{\ell=\mathrm{policy}}$ with $\ell\sim c_{\phi}(\cdot\mid a,q)$ is an unbiased estimator of the reward $r_{\phi}(a,q)=c_{\phi}(\ell=\mathrm{expert}\mid a,q)-c_{\phi}(\ell=\mathrm{policy}\mid a,q)$.
Proof.
We compute the expectation of $\hat{r}_{\phi}(a,q)$ over the sampling distribution $c_{\phi}(\cdot\mid a,q)$:
$\mathbb{E}_{\ell\sim c_{\phi}(\cdot\mid a,q)}[\hat{r}_{\phi}(a,q)]$ $=c_{\phi}(\ell=\mathrm{expert}\mid a,q)\cdot(1)+c_{\phi}(\ell=\mathrm{policy}\mid a,q)\cdot(-1)$
$=c_{\phi}(\ell=\mathrm{expert}\mid a,q)-c_{\phi}(\ell=\mathrm{policy}\mid a,q).$
Since $c_{\phi}(\ell=\mathrm{policy}\mid a,q)=1-c_{\phi}(\ell=\mathrm{expert}\mid a,q)$, we have:
$\mathbb{E}[\hat{r}_{\phi}(a,q)]=c_{\phi}(\ell=\mathrm{expert}\mid a,q)-c_{\phi}(\ell=\mathrm{policy}\mid a,q)=r_{\phi}(a,q).$

### v3

@@ 3.4 RARO: Relativistic Adversarial Reasoning Optimization
Finally, we identify the key techniques for stable and efficient training. First, we use a shared LLM for both the critic and the policy, which reduces memory usage and promotes generalization. This allows us to employ data mixing, where policy and critic rollouts are combined in a single batch, simplifying the training loop. To prevent the critic from catastrophic forgetting, we sample critic prompts from a replay buffer that stores all past expert and policy answers. Finally, we incorporate several practical improvements to the GRPO algorithm, such as over-length filtering and removing advantage/length normalization. For full implementation details, please refer to Appendix C.1.
@@
Inputs: Dataset $D=\{(q_{i},a_{i})\}$; Tie reward $\tau_{\mathrm{pol}},\tau_{\mathrm{crit}}$; Loss weight $\lambda_{\mathrm{pol}},\lambda_{\mathrm{crit}}$; Batch $B$; Rollout $K$.
@@ 4.1 Tasks & Datasets
Then, we evaluate our method on the domain of general math reasoning problems using the DeepMath dataset (he2025deepmath103klargescalechallengingdecontaminated). Compared to Countdown, answer verification in the general math domain is significantly more challenging, often requiring solving the problem from scratch. Through this task, we aim to stress test our method on difficult reasoning environments where verification is as difficult as generation.
@@
Finally, we extend our method to its intended setting of non-verifiable, open-ended reasoning tasks. While there exists benchmarks for non-verifiable domains (arora2025healthbenchevaluatinglargelanguage; paech2024eqbenchemotionalintelligencebenchmark), there is no official training data associated with them. Thus, we choose to evaluate our method and baselines on a custom Poetry Writing dataset. Unlike the math tasks, poetry writing does not admit an objective verifier. Thus, for evaluation, we use GPT-5  (openai2025gpt5) as a judge to evaluate poems in both isolation and in comparison to the expert poem (see Appendix C.2 for details). This task represents the non-verifiable regime that our method aims to handle, where explicit reasoning could significantly improve quality.
@@ 4.2 Baselines
The SFT baseline trains the base models to directly maximize the conditional log-likelihood of the expert answer given the question, representing the standard use of demonstration data.
@@ 4.3 Training & Evaluation Setup
We evaluate our method and baselines on the Qwen2.5 (qwen2025qwen25technicalreport) family of models, and to focus on improving reasoning performance rather than language understanding, we initialize from the instruction-tuned checkpoints instead of the pretrained model checkpoints. We select the Qwen2.5 family because they are popular non-reasoning LLMs, allowing us to study the effectiveness of our method on eliciting reasoning behaviors in a controlled manner.
Countdown and DeepMath are evaluated with a ground-truth verifier, while Poetry Writing is evaluated with GPT-5 as a judge in two fashions: a scalar score normalized from 1-7 to 0-100 and a win-rate against the expert poem. For both fashions, we prompt GPT-5 to focus its evaluations on prompt adherence and literary qualities. See Appendix C.2 for further details.
Each dataset is split into train, validation, and test sets, and for each method, we select the non-initial checkpoint with the highest validation performance. We match dataset splits, rollout budgets, hyperparameters, and sampling configurations for all methods when possible to ensure a fair comparison. Unless otherwise specified, all methods are trained and evaluated with a reasoning budget of 2048 tokens. Full implementation details and hyperparameters are provided in Appendix C.
5 Main Results
@@ 5.1 Countdown
5.2 DeepMath
Next, we evaluate RARO on the DeepMath dataset, a collection of general math problems. For the DeepMath task, we focus on scaling our method and baselines with respect to model size instead of reasoning budget, as it is a much more difficult setting where model capacity is a real bottleneck in performance.
@@ 5.2 DeepMath
accuracy ($\%$) $\uparrow$ score (0-100) $\uparrow$
win-rate ($\%$) $\uparrow$
(vs. expert)
@@
As reported in Table 2, RARO consistently outperforms all verifier-free baselines across model scales. With the 1.5B model, we achieve $41.3\%$ accuracy compared to $37.7\%$ for the best baseline (RL-Logit), an improvement of $3.6\%$. This advantage grows with model size: at 3B, our method ($49.1\%$) surpasses the best baseline (RL-Logit, $43.1\%$) by $6.0\%$, and at 7B, it reaches $57.5\%$, beating it by $8.2\%$. These results demonstrate that our adversarial learning framework provides a strong signal for reasoning that outperforms not only purely supervised approaches like SFT or Rationalization but also RL-Logit.
@@
We further analyze the training dynamics of RARO on DeepMath. As shown in Figure 4 and further in Appendix B, our coupled training objective maintains a robust equilibrium, allowing the policy to steadily improve its reasoning capabilities and response length without collapsing. This stability confirms the robustness of our optimization procedure.
@@
Another key advantage of RARO is that our learned critic enables natural Test-Time Scaling (TTS) to further improve the policy’s performance. Specifically, our critic’s pairwise comparison setup allows for a single-elimination tournament with the critic as the judge (see Algorithm 3), enabling further policy improvements with additional rollouts. As shown in Figure 5 (and detailed in Table 9), increasing the number of rollouts from 1 to 16 consistently improves performance. Notably, with 16 rollouts, RARO achieves $53.6\%$ on the 1.5B model and $57.9\%$ on the 3B model. When comparing against the RLVR baseline with the same TTS strategy (Table 9), we observe that RARO achieves a similar rate of improvement. This result highlights that RARO, when combined with test-time search, can scale effectively, matching the scaling trends of models trained with oracle verifiers.
5.3 Poetry Writing
@@
Table 12 reveals a striking performance gap between RARO and baselines. While SFT and Rationalization achieve modest win-rates against expert poetry (peaking at $5.9\%$ with the 7B model), RARO reaches $25.0\%$, a four-fold improvement. This advantage is also reflected in the scoring evaluation, where RARO consistently surpasses baselines (e.g., $67.8$ vs. $53.7$ for SFT at 1.5B). Notably, RL-Logit, leading baseline for DeepMath, fails to produce competitive results, yielding near-zero improvement over the base model ($36.4$ vs. $35.0$ at 1.5B) for both the win-rate and scoring evaluation.
@@
A key result is the scalability of RARO with model size in the creative domain. As we increase model capacity from 1.5B to 7B, the win-rate against expert human poems grows substantially, from $7.8\%$ to $25.0\%$. The scoring evaluation similarly improves from $67.8$ to $77.3$. This trend shows that just like verifiable domains, RARO continues to effectively scale with model size in open-ended domains.
Countdown Use the four numbers 19, 11, 48, and 6 to create an expression that equals 24.
RARO Policy To find an expression using the four numbers 19, 11, 48, and 6 that equals 24, let’s try different combinations. 1. (48 - 11) * (19 - 6) = 37 * 13 = 481 (too high) 2. (48 - 19) * (11 - 6) = 29 * 5 = 145 (too high) 20. 48 / (19 - (11 + 6)) = 48 / (19 - 17) = 48 / 2 = 24 (correct) Answer: 48/(19(11+6))
DeepMath Find the equation of the normal line to the curve $(x^{2}+y^{2})^{3/2}=2xy$ at the point $\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$.
RARO Policy To find the equation of the normal line to the curve $(x^{2}+y^{2})^{3/2}=2xy$ at the point $\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$, we need to follow these steps: 1. Differentiate the curve implicitly to find the slope of the tangent line at the given point. 2. Use the slope of the tangent line to find the slope of the normal line. 3. Write the equation of the normal line using the point-slope form. Let’s start with… Answer: y=x
Poetry Writing a poem about a child’s fragmented memory of a father leaving, told in intimate, sensory free verse
RARO Policy To create such a poem, I’d: 1. Focus on sensory details that might linger in a child’s memory. 4. Incorporate physical and emotional sensations that capture the essence of the scene. Answer:
One Memory of My Father Leaving
There was the sound of a suitcase rolling on the hallway carpet; or a man thinking and trying not to cry.
Figure 6: Qualitative examples for Countdown (1.5B), DeepMath (7B), and Poetry Writing (7B). For each task, we show the input (top) and the truncated reasoning and answer sampled from policies trained with RARO (bottom). See Appendix E for full examples as well as example critic inputs and outputs.
@@
Qualitatively, RARO induces explicit planning and reasoning behaviors. As shown in Figure 6, the model learns to decompose complex problems into structured plans. Specifically, in Poetry Writing, it decomposes the prompt into key themes (e.g., “sensory details”, “fragmented memory”) and stylistic constraints (e.g., “free verse”, “unrhymed, natural cadences”) before generating the poem. This demonstrates that RARO effectively elicits reasonings that align the model’s output to the task requirements.
@@ 6 Conclusion & Future Work
We introduced RARO (Relativistic Adversarial Reasoning Optimization), a novel approach to training reasoning LLMs using only expert demonstrations, thereby bypassing the need for task-specific verifiers or expensive preference annotations. By formulating the problem as Inverse Reinforcement Learning and incorporating a relativistic critic setup, we obtain a principled and stable adversarial training algorithm that yields strong reasoning capabilities.
Our experiments demonstrate the effectiveness of RARO: (i) on the controlled Countdown task, it not only outperforms verifier-free baselines but also nearly matches the performance of RLVR; (ii) on the general math domain, it exhibits similar desirable scalability trends to RLVR while outperforming baselines without verification; and (iii) on the open-ended Poetry Writing task, it successfully elicits emerging specialized reasoning capabilities and significantly surpasses all baselines. Together, these findings suggest that RARO is a promising and practical approach for training reasoning models without reliance on explicit verifiers.
@@ A.2 Proof of Reward Gradient
Using the closed-form policy, the gradient of the maximum likelihood objective $\mathcal{L}(\phi)=\mathbb{E}_{(q,a)\sim\hat{p}_{D}}\!\left[\log\pi_{\theta^{\star}(\phi)}(a\mid q)\right]$ with respect to $\phi$ is:
@@ A.3 Derivation of Reasoning Reward Gradient
Let the reward be parameterized by a critic $c_{\phi}(\ell\mid a,q)$ with labels $\ell\in\{\mathrm{expert},\mathrm{policy}\}$ as $r_{\phi}(a,q)=c_{\phi}(\ell=\mathrm{expert}\mid a,q)$. The gradient of the loss with respect to critic parameters $\phi$ is:
@@
We parameterize the reward using the probability of the answer being expert-generated, as predicted by a binary classifier (critic) $c_{\phi}(\ell\mid a,q)$ where $\ell\in\{\mathrm{expert},\mathrm{policy}\}$:
$r_{\phi}(a,q)=c_{\phi}(\ell=\mathrm{expert}\mid a,q).$
@@
$\mathbb{E}_{\ell\sim c_{\phi}(\cdot\mid a,q)}[\mathbbm{1}_{\ell=\mathrm{expert}}\nabla_{\phi}\log c_{\phi}(\ell\mid a,q)]$ $=\mathbbm{1}_{\mathrm{expert}=\mathrm{expert}}p_{E}\nabla_{\phi}\log p_{E}+\mathbbm{1}_{\mathrm{policy}=\mathrm{expert}}p_{P}\nabla_{\phi}\log p_{P}$
$=1\cdot\nabla_{\phi}p_{E}+0$
$=\nabla_{\phi}p_{E}.$
@@
$\nabla_{\phi}r_{\phi}(a,q)=\mathbb{E}_{\ell\sim c_{\phi}(\cdot\mid a,q)}\Big[\mathbbm{1}_{\ell=\mathrm{expert}}\nabla_{\phi}\log c_{\phi}(\ell\mid a,q)\Big].$
@@
1. Expert Term ($(q,a)\sim\hat{p}_{D}$): The answers come from the expert distribution, so the correct label is $\mathrm{expert}$.
$\mathbb{E}_{a\sim\hat{p}_{a\mid q}}[\nabla_{\phi}r_{\phi}(a,q)]$ $=\mathbb{E}_{a\sim\hat{p}_{a\mid q}}\Big[\mathbb{E}_{\ell\sim c_{\phi}}\Big[\mathbbm{1}_{\ell=\mathrm{expert}}\nabla_{\phi}\log c_{\phi}(\ell\mid a,q)\Big]\Big].$
This corresponds to a reward signal of +1+1 when $\ell=\mathrm{expert}$ (correct) and 0 when $\ell=\mathrm{policy}$ (incorrect).
@@
$-\mathbb{E}_{a\sim\pi_{\theta}}[\nabla_{\phi}r_{\phi}(a,q)]$ $=-\mathbb{E}_{a\sim\pi_{\theta}}[\nabla_{\phi}p_{E}]$
$=\mathbb{E}_{a\sim\pi_{\theta}}[\nabla_{\phi}p_{P}]\quad(\text{since }\nabla_{\phi}p_{E}+\nabla_{\phi}p_{P}=0)$
$=\mathbb{E}_{a\sim\pi_{\theta}}\Big[\mathbb{E}_{\ell\sim c_{\phi}}\Big[\mathbbm{1}_{\ell=\mathrm{policy}}\nabla_{\phi}\log c_{\phi}(\ell\mid a,q)\Big]\Big].$
Here, the answers come from the policy, so the correct label is $\mathrm{policy}$. This corresponds to a reward signal of +1+1 when $\ell=\mathrm{policy}$ (correct) and 0 when $\ell=\mathrm{expert}$ (incorrect).